# BOP Challenge 2023 on Detection, Segmentation and Pose Estimation of Seen and Unseen Rigid Objects

### v1

@@ 4.1 Experimental setup
| 34 | ZebraPoseSAT-EffnetB4 (PBR_Only) [54] | 2022 | DNN | Object | FCOS | – | RGB | PBR | RGB | 72.1 | 72.3 | 71.7 | 54.5 | 41.0 | 88.2 | 69.1 | 67.0 |
| 35 | Extended_FCOS+PFA-PBR-RGB [28] | 2022 | DNN | Dataset | Extended FCOS | PFA | RGB | PBR | RGB | 74.5 | 71.9 | 73.2 | 60.0 | 35.3 | 84.1 | 64.8 | 66.3 |
3.50 |
| 36 | PFA-cosypose [28, 33] | 2022 | DNN | Dataset | MaskRCNN | PFA | RGB-D | PBR+real | RGB | 67.4 | 73.8 | 83.7 | 59.6 | 24.6 | 71.2 | 80.7 | 65.9 |
@@
| 38 | SurfEmb-PBR-RGB [16] | 2022 | DNN | Dataset | Default (PBR) | Custom | RGB | PBR | RGB | 66.3 | 73.5 | 71.5 | 58.8 | 41.3 | 79.1 | 64.7 | 65.0 |
8.89 |
| 39 | Koenig-Hybrid-DL-PointPairs [32] | 2020 | DNN/PPF | Dataset | Retina/MaskRCNN | ICP | RGB | Synt+real | RGB-D | 63.1 | 65.5 | 92.0 | 43.0 | 48.3 | 65.1 | 70.1 | 63.9 |
0.63 |
| 40 | CosyPose-ECCV20-SYNT+REAL-1VIEW [33] | 2020 | DNN | Dataset | Default (synt+real) | DeepIM | RGB | PBR+real | RGB | 63.3 | 72.8 | 82.3 | 58.3 | 21.6 | 65.6 | 82.1 | 63.7 |
0.45 |
| 41 | CRT-6D | 2022 | DNN | Dataset | Default (synt+real) | Custom | RGB | PBR+real | RGB | 66.0 | 64.4 | 78.9 | 53.7 | 20.8 | 60.3 | 75.2 | 59.9 |
0.06 |
| 42 | Pix2Pose-BOP20_w/ICP-ICCV19 [47] | 2020 | DNN | Object | MaskRCNN | ICP | RGB | PBR+real | RGB-D | 58.8 | 51.2 | 82.0 | 39.0 | 35.1 | 69.5 | 78.0 | 59.1 |
4.84 |
| 43 | ZTE_PPF | 2022 | DNN/PPF | Dataset | Default (synt+real) | ICP | RGB | PBR+real | RGB-D | 66.3 | 37.4 | 90.4 | 39.6 | 47.0 | 73.5 | 50.2 | 57.8 |
0.90 |
| 44 | CosyPose-ECCV20-PBR-1VIEW [33] | 2020 | DNN | Dataset | Default (PBR) | DeepIM | RGB | PBR | RGB | 63.3 | 64.0 | 68.5 | 58.3 | 21.6 | 65.6 | 57.4 | 57.0 |
0.48 |
| 45 | Vidal-Sensors18 [61] | 2019 | PPF | – | – | ICP | – | – | D | 58.2 | 53.8 | 87.6 | 39.3 | 43.5 | 70.6 | 45.0 | 56.9 |
3.22 |
| 46 | CDPNv2_BOP20 (RGB-only & ICP) [35] | 2020 | DNN | Object | FCOS | ICP | RGB | Synt+real | RGB-D | 63.0 | 46.4 | 91.3 | 45.0 | 18.6 | 71.2 | 61.9 | 56.8 |
1.46 |
@@
| 48 | MRPE-PBR-RGB-SModel | 2023 | DNN | Dataset | Default GDRNPPDet | – | RGB | PBR | RGB | 71.5 | 72.9 | 20.6 | 46.2 | 35.3 | 76.5 | 55.2 | 54.0 |
0.10 |
| 49 | CDPNv2_BOP20 (PBR-only & ICP) [35] | 2020 | DNN | Object | FCOS | ICP | RGB | PBR | RGB-D | 63.0 | 43.5 | 79.1 | 45.0 | 18.6 | 71.2 | 53.2 | 53.4 |
1.49 |
| 50 | CDPNv2_BOP20 (RGB-only) [35] | 2020 | DNN | Object | FCOS | – | RGB | Synt+real | RGB | 62.4 | 47.8 | 77.2 | 47.3 | 10.2 | 72.2 | 53.2 | 52.9 |
0.94 |
@@
| 52 | Drost-CVPR10-3D-Only [11] | 2019 | PPF | – | – | ICP | – | – | D | 52.7 | 44.4 | 77.5 | 38.8 | 31.6 | 61.5 | 34.4 | 48.7 |
7.70 |
| 53 | CDPN_BOP19 (RGB-only) [35] | 2020 | DNN | Object | RetinaNet | – | RGB | Synt+real | RGB | 56.9 | 49.0 | 76.9 | 32.7 |
6.7 |
67.2 | 45.7 | 47.9 |
0.48 |
| 54 | CDPNv2_BOP20 (PBR-only & RGB-only) [35] | 2020 | DNN | Object | FCOS | – | RGB | PBR | RGB | 62.4 | 40.7 | 58.8 | 47.3 | 10.2 | 72.2 | 39.0 | 47.2 |
0.98 |
| 55 | leaping from 2D to 6D [39] | 2020 | DNN | Object | Unknown | – | RGB | Synt+real | RGB | 52.5 | 40.3 | 75.1 | 34.2 |
7.7 |
65.8 | 54.3 | 47.1 |
0.43 |
| 56 | EPOS-BOP20-PBR [21] | 2020 | DNN | Dataset | – | – | RGB | PBR | RGB | 54.7 | 46.7 | 55.8 | 36.3 | 18.6 | 58.0 | 49.9 | 45.7 |
1.87 |
| 57 | Drost-CVPR10-3D-Only-Faster [11] | 2019 | PPF | – | – | ICP | – | – | D | 49.2 | 40.5 | 69.6 | 37.7 | 27.4 | 60.3 | 33.0 | 45.4 |
1.38 |
| 58 | Félix&Neves-ICRA2017-IET2019 [52, 49] | 2019 | DNN/PPF | Dataset | MaskRCNN | ICP | RGB-D | Synt+real | RGB-D | 39.4 | 21.2 | 85.1 | 32.3 |
6.9 |
52.9 | 51.0 | 41.2 | 55.78 |
| 59 | Sundermeyer-IJCV19+ICP [57] | 2019 | DNN | Object | RetinaNet | ICP | RGB | Synt+real | RGB-D | 23.7 | 48.7 | 61.4 | 28.1 | 15.8 | 50.6 | 50.5 | 39.8 |
0.86 |
| 60 | Zhigang-CDPN-ICCV19 [35] | 2019 | DNN | Object | RetinaNet | – | RGB | Synt+real | RGB | 37.4 | 12.4 | 75.7 | 25.7 |
7.0 |
47.0 | 42.2 | 35.3 |
0.51 |
| 61 | PointVoteNet2 [13] | 2020 | DNN | Object | – | ICP | RGB-D | PBR | RGB-D | 65.3 |
0.4 |
67.3 | 26.4 |
0.1 |
55.6 | 30.8 | 35.1 |
| 62 | Pix2Pose-BOP20-ICCV19 [47] | 2020 | DNN | Object | MaskRCNN | – | RGB | PBR+real | RGB | 36.3 | 34.4 | 42.0 | 22.6 | 13.4 | 44.6 | 45.7 | 34.2 |
1.22 |
| 63 | Sundermeyer-IJCV19[57] | 2019 | DNN | Object | RetinaNet | – | RGB | Synt+real | RGB | 14.6 | 30.4 | 40.1 | 21.7 | 10.1 | 34.6 | 44.6 | 28.0 |
0.20 |
| 64 | SingleMultiPathEncoder-CVPR20 [55] | 2020 | DNN | All | MaskRCNN | – | RGB | Synt+real | RGB | 21.7 | 31.0 | 33.4 | 17.5 |
6.7 |
29.3 | 28.9 | 24.1 |
0.19 |
| 65 | DPOD (synthetic) [67] | 2019 | DNN | Dataset | – | – | RGB | Synt | RGB | 16.9 |
8.1 |
24.2 | 13.0 |
0.0 |
28.6 | 22.2 | 16.1 |
0.23 |
| # | Method | Year | Type | DNN per | Det./seg. | Refinement | Train im. | …type | Test im. |
LM-O |
T-LESS |
TUD-L |
IC-BIN |
ITODD |
HB |
YCB-V |
Time |
| 1 | GenFlow-MultiHypo16 [41] | 2023 | DNN | All | CNOS-fastSAM | Recurrent Flow | RGB-D | PBR | RGB-D | 63.5 | 52.1 | 86.2 | 53.4 | 55.4 | 77.9 | 83.3 | 67.4 |
34.58 |
| 2 | GenFlow-MultiHypo [41] | 2023 | DNN | All | CNOS-fastSAM | Recurrent Flow | RGB-D | PBR | RGB-D | 62.2 | 50.9 | 84.9 | 52.4 | 54.4 | 77.0 | 81.8 | 66.2 |
21.46 |
@@
| 5 | SAM6D-CNOSmask [36, 44] | 2023 | DNN | All | CNOS-fastSAM | Cross-attention | RGB-D | PBR | RGB-D | 64.8 | 48.3 | 79.4 | 50.4 | 35.1 | 72.7 | 80.4 | 61.6 |
3.87 |
@@
| 7 | ZeroPose-Multi-Hypo-Refinement [4, 44] | 2023 | DNN | All | CNOS-fastSAM | MegaPose | RGB-D | PBR+real | RGB-D | 53.8 | 40.0 | 83.5 | 39.2 | 52.1 | 65.3 | 65.3 | 57.0 |
16.17 |
| 8 | GenFlow-MultiHypo-RGB | 2023 | DNN | All | CNOS-fastSAM | Recurrent Flow | RGB-D | PBR | RGB | 56.3 | 52.3 | 68.4 | 45.3 | 39.5 | 73.9 | 63.3 | 57.0 |
20.89 |
| 9 | Megapose-CNOS_fastSAM+Multih-10 [34, 44] | 2023 | DNN | All | CNOS-fastSAM | MegaPose | RGB | PBR | RGB | 56.0 | 50.8 | 68.7 | 41.9 | 34.6 | 70.6 | 62.0 | 54.9 |
53.88 |
| 10 | Megapose-CNOS_fastSAM+Multih [34, 44] | 2023 | DNN | All | CNOS-fastSAM | MegaPose | RGB | PBR | RGB | 56.0 | 50.7 | 68.4 | 41.4 | 33.8 | 70.4 | 62.1 | 54.7 |
47.39 |
| 11 | ZeroPose-Multi-Hypo-Refinement [4, 44] | 2023 | DNN | All | SAM + ImageBind | MegaPose | RGB-D | PBR+real | RGB-D | 49.3 | 34.2 | 79.0 | 39.6 | 46.5 | 62.9 | 62.3 | 53.4 |
18.97 |
| 12 | MegaPose-CNOS_fastSAM [34, 44] | 2023 | DNN | All | CNOS-fastSAM | MegaPose | RGB | PBR | RGB | 49.9 | 47.7 | 65.3 | 36.7 | 31.5 | 65.4 | 60.1 | 50.9 |
31.72 |
| 13 | ZeroPose-One-Hypo [4] | 2023 | DNN | All | SAM + ImageBind | MegaPose | RGB-D | PBR+real | RGB-D | 27.2 | 15.6 | 53.6 | 30.7 | 36.2 | 46.2 | 34.1 | 34.8 |
9.76 |
| 14 | GenFlow-coarse | 2023 | DNN | All | CNOS-fastSAM | – | RGB | PBR | RGB | 25.0 | 21.5 | 30.0 | 16.8 | 15.4 | 28.3 | 27.7 | 23.5 |
3.84 |
Method |
…based on |
Year | Data | …type | AP | Time |
| 1 |
GDet2023 |
YOLOv8 |
2023 | RGB | PBR+real | 79.8 | .204 |
| 2 |
GDRNPPDet |
YOLOX |
2022 | RGB | PBR+real | 77.3 | .081 |
| 3 |
GDet2023-PBR |
YOLOv8 |
2023 | RGB | PBR | 76.9 | .204 |
| 4 |
GDRNPPDet |
YOLOX |
2022 | RGB | PBR | 73.8 | .081 |
| 5 |
Extended_FCOS |
FCOS |
2022 | RGB | PBR+real | 72.1 | .030 |
| 6 |
Extended_FCOS |
FCOS |
2022 | RGB | PBR | 66.7 | .030 |
| 7 |
DLZDet |
DLZDet |
2022 | RGB | PBR | 65.6 | - |
| 8 |
CosyPose |
Mask R-CNN |
2020 | RGB | PBR+real | 60.5 | .054 |
| 9 |
CosyPose |
Mask R-CNN |
2020 | RGB | PBR | 55.7 | .055 |
| 10 |
FCOS-CDPN |
FCOS |
2022 | RGB | PBR | 50.7 | .047 |
Method |
…based on |
Year | Data | …type | AP | Time |
| 1 |
ZebraPoseSAT |
GDRNPP+Zebra |
2023 | RGB | PBR+real | 61.9 | .080 |
| 2 |
ZebraPoseSAT |
GDRNPP+Zebra |
2022 | RGB | PBR+real | 58.7 | .080 |
| 3 |
ZebraPoseSAT |
CDPNv2+Zebra |
2023 | RGB | PBR+real | 57.9 | .080 |
| 4 |
ZebraPoseSAT |
CDPNv2+Zebra |
2022 | RGB | PBR+real | 57.8 | .080 |
| 5 |
ZebraPoseSAT |
CosyPose+Zebra |
2022 | RGB | PBR | 53.8 | .080 |
| 6 |
ZebraPoseSAT |
CDPNv2+Zebra |
2022 | RGB | PBR | 52.3 | .080 |
| 7 |
DLZDet |
DLZDet |
2022 | RGB | PBR+real | 49.6 | - |
| 8 |
DLZDet |
DLZDet |
2022 | RGB | PBR | 42.9 | - |
| 9 |
CosyPose |
Mask R-CNN |
2020 | RGB | PBR+real | 40.5 | .054 |
| 10 |
CosyPose |
Mask R-CNN |
2020 | RGB | PBR | 36.2 | .055 |
Method |
Year | Train. im. | …type | Test. im. | AP | Time |
| 1 |
CNOS_FastSAM [44] |
2023 | RGB | PBR | RGB | 42.8 | 0.221 |
| 2 |
CNOS_SAM [44] |
2023 | RGB | PBR | RGB | 36.1 | 1.847 |
| 3 |
ZeroPose [4] |
2023 | RGB | PBR | RGB | 34.1 | 3.821 |

### v2

@@ 4.1 Experimental setup
| 34 | ZebraPoseSAT-EffnetB4 (PBR_Only) [53] | 2022 | DNN | Object | FCOS | – | RGB | PBR | RGB | 72.1 | 72.3 | 71.7 | 54.5 | 41.0 | 88.2 | 69.1 | 67.0 | – |
| 35 | Extended_FCOS+PFA-PBR-RGB [27] | 2022 | DNN | Dataset | Extended FCOS | PFA | RGB | PBR | RGB | 74.5 | 71.9 | 73.2 | 60.0 | 35.3 | 84.1 | 64.8 | 66.3 | 3.50 |
| 36 | PFA-cosypose [27, 32] | 2022 | DNN | Dataset | MaskRCNN | PFA | RGB-D | PBR+real | RGB | 67.4 | 73.8 | 83.7 | 59.6 | 24.6 | 71.2 | 80.7 | 65.9 | – |
@@
| 38 | SurfEmb-PBR-RGB [15] | 2022 | DNN | Dataset | Default (PBR) | Custom | RGB | PBR | RGB | 66.3 | 73.5 | 71.5 | 58.8 | 41.3 | 79.1 | 64.7 | 65.0 | 8.89 |
| 39 | Koenig-Hybrid-DL-PointPairs [31] | 2020 | DNN/PPF | Dataset | Retina/MaskRCNN | ICP | RGB | Synt+real | RGB-D | 63.1 | 65.5 | 92.0 | 43.0 | 48.3 | 65.1 | 70.1 | 63.9 | 0.63 |
| 40 | CosyPose-ECCV20-SYNT+REAL-1VIEW [32] | 2020 | DNN | Dataset | Default (synt+real) | DeepIM | RGB | PBR+real | RGB | 63.3 | 72.8 | 82.3 | 58.3 | 21.6 | 65.6 | 82.1 | 63.7 | 0.45 |
| 41 | CRT-6D | 2022 | DNN | Dataset | Default (synt+real) | Custom | RGB | PBR+real | RGB | 66.0 | 64.4 | 78.9 | 53.7 | 20.8 | 60.3 | 75.2 | 59.9 | 0.06 |
| 42 | Pix2Pose-BOP20_w/ICP-ICCV19 [46] | 2020 | DNN | Object | MaskRCNN | ICP | RGB | PBR+real | RGB-D | 58.8 | 51.2 | 82.0 | 39.0 | 35.1 | 69.5 | 78.0 | 59.1 | 4.84 |
| 43 | ZTE_PPF | 2022 | DNN/PPF | Dataset | Default (synt+real) | ICP | RGB | PBR+real | RGB-D | 66.3 | 37.4 | 90.4 | 39.6 | 47.0 | 73.5 | 50.2 | 57.8 | 0.90 |
| 44 | CosyPose-ECCV20-PBR-1VIEW [32] | 2020 | DNN | Dataset | Default (PBR) | DeepIM | RGB | PBR | RGB | 63.3 | 64.0 | 68.5 | 58.3 | 21.6 | 65.6 | 57.4 | 57.0 | 0.48 |
| 45 | Vidal-Sensors18 [60] | 2019 | PPF | – | – | ICP | – | – | D | 58.2 | 53.8 | 87.6 | 39.3 | 43.5 | 70.6 | 45.0 | 56.9 | 3.22 |
| 46 | CDPNv2_BOP20 (RGB-only & ICP) [34] | 2020 | DNN | Object | FCOS | ICP | RGB | Synt+real | RGB-D | 63.0 | 46.4 | 91.3 | 45.0 | 18.6 | 71.2 | 61.9 | 56.8 | 1.46 |
@@
| 48 | MRPE-PBR-RGB-SModel | 2023 | DNN | Dataset | Default GDRNPPDet | – | RGB | PBR | RGB | 71.5 | 72.9 | 20.6 | 46.2 | 35.3 | 76.5 | 55.2 | 54.0 | 0.10 |
| 49 | CDPNv2_BOP20 (PBR-only & ICP) [34] | 2020 | DNN | Object | FCOS | ICP | RGB | PBR | RGB-D | 63.0 | 43.5 | 79.1 | 45.0 | 18.6 | 71.2 | 53.2 | 53.4 | 1.49 |
| 50 | CDPNv2_BOP20 (RGB-only) [34] | 2020 | DNN | Object | FCOS | – | RGB | Synt+real | RGB | 62.4 | 47.8 | 77.2 | 47.3 | 10.2 | 72.2 | 53.2 | 52.9 | 0.94 |
@@
| 52 | Drost-CVPR10-3D-Only [10] | 2019 | PPF | – | – | ICP | – | – | D | 52.7 | 44.4 | 77.5 | 38.8 | 31.6 | 61.5 | 34.4 | 48.7 | 7.70 |
| 53 | CDPN_BOP19 (RGB-only) [34] | 2020 | DNN | Object | RetinaNet | – | RGB | Synt+real | RGB | 56.9 | 49.0 | 76.9 | 32.7 | 6.7 | 67.2 | 45.7 | 47.9 | 0.48 |
| 54 | CDPNv2_BOP20 (PBR-only & RGB-only) [34] | 2020 | DNN | Object | FCOS | – | RGB | PBR | RGB | 62.4 | 40.7 | 58.8 | 47.3 | 10.2 | 72.2 | 39.0 | 47.2 | 0.98 |
| 55 | leaping from 2D to 6D [38] | 2020 | DNN | Object | Unknown | – | RGB | Synt+real | RGB | 52.5 | 40.3 | 75.1 | 34.2 | 7.7 | 65.8 | 54.3 | 47.1 | 0.43 |
| 56 | EPOS-BOP20-PBR [20] | 2020 | DNN | Dataset | – | – | RGB | PBR | RGB | 54.7 | 46.7 | 55.8 | 36.3 | 18.6 | 58.0 | 49.9 | 45.7 | 1.87 |
| 57 | Drost-CVPR10-3D-Only-Faster [10] | 2019 | PPF | – | – | ICP | – | – | D | 49.2 | 40.5 | 69.6 | 37.7 | 27.4 | 60.3 | 33.0 | 45.4 | 1.38 |
| 58 | Félix&Neves-ICRA2017-IET2019 [51, 48] | 2019 | DNN/PPF | Dataset | MaskRCNN | ICP | RGB-D | Synt+real | RGB-D | 39.4 | 21.2 | 85.1 | 32.3 | 6.9 | 52.9 | 51.0 | 41.2 | 55.78 |
| 59 | Sundermeyer-IJCV19+ICP [56] | 2019 | DNN | Object | RetinaNet | ICP | RGB | Synt+real | RGB-D | 23.7 | 48.7 | 61.4 | 28.1 | 15.8 | 50.6 | 50.5 | 39.8 | 0.86 |
| 60 | Zhigang-CDPN-ICCV19 [34] | 2019 | DNN | Object | RetinaNet | – | RGB | Synt+real | RGB | 37.4 | 12.4 | 75.7 | 25.7 | 7.0 | 47.0 | 42.2 | 35.3 | 0.51 |
| 61 | PointVoteNet2 [12] | 2020 | DNN | Object | – | ICP | RGB-D | PBR | RGB-D | 65.3 | 0.4 | 67.3 | 26.4 | 0.1 | 55.6 | 30.8 | 35.1 | – |
| 62 | Pix2Pose-BOP20-ICCV19 [46] | 2020 | DNN | Object | MaskRCNN | – | RGB | PBR+real | RGB | 36.3 | 34.4 | 42.0 | 22.6 | 13.4 | 44.6 | 45.7 | 34.2 | 1.22 |
| 63 | Sundermeyer-IJCV19[56] | 2019 | DNN | Object | RetinaNet | – | RGB | Synt+real | RGB | 14.6 | 30.4 | 40.1 | 21.7 | 10.1 | 34.6 | 44.6 | 28.0 | 0.20 |
| 64 | SingleMultiPathEncoder-CVPR20 [54] | 2020 | DNN | All | MaskRCNN | – | RGB | Synt+real | RGB | 21.7 | 31.0 | 33.4 | 17.5 | 6.7 | 29.3 | 28.9 | 24.1 | 0.19 |
| 65 | DPOD (synthetic) [66] | 2019 | DNN | Dataset | – | – | RGB | Synt | RGB | 16.9 | 8.1 | 24.2 | 13.0 | 0.0 | 28.6 | 22.2 | 16.1 | 0.23 |
| # | Method | Year | Type | DNN per | Det./seg. | Refinement | Train im. | …type | Test im. | LM-O | T-LESS | TUD-L | IC-BIN | ITODD | HB | YCB-V | Time | |
| 1 | GenFlow-MultiHypo16 [40] | 2023 | DNN | All | CNOS-fastSAM | Recurrent Flow | RGB-D | PBR | RGB-D | 63.5 | 52.1 | 86.2 | 53.4 | 55.4 | 77.9 | 83.3 | 67.4 | 34.58 |
| 2 | GenFlow-MultiHypo [40] | 2023 | DNN | All | CNOS-fastSAM | Recurrent Flow | RGB-D | PBR | RGB-D | 62.2 | 50.9 | 84.9 | 52.4 | 54.4 | 77.0 | 81.8 | 66.2 | 21.46 |
@@
| 5 | SAM6D-CNOSmask [35, 43] | 2023 | DNN | All | CNOS-fastSAM | Cross-attention | RGB-D | PBR | RGB-D | 64.8 | 48.3 | 79.4 | 50.4 | 35.1 | 72.7 | 80.4 | 61.6 | 3.87 |
@@
| 7 | ZeroPose-Multi-Hypo-Refinement [3, 43] | 2023 | DNN | All | CNOS-fastSAM | MegaPose | RGB-D | PBR+real | RGB-D | 53.8 | 40.0 | 83.5 | 39.2 | 52.1 | 65.3 | 65.3 | 57.0 | 16.17 |
| 8 | GenFlow-MultiHypo-RGB | 2023 | DNN | All | CNOS-fastSAM | Recurrent Flow | RGB-D | PBR | RGB | 56.3 | 52.3 | 68.4 | 45.3 | 39.5 | 73.9 | 63.3 | 57.0 | 20.89 |
| 9 | Megapose-CNOS_fastSAM+Multih-10 [33, 43] | 2023 | DNN | All | CNOS-fastSAM | MegaPose | RGB | PBR | RGB | 56.0 | 50.8 | 68.7 | 41.9 | 34.6 | 70.6 | 62.0 | 54.9 | 53.88 |
| 10 | Megapose-CNOS_fastSAM+Multih [33, 43] | 2023 | DNN | All | CNOS-fastSAM | MegaPose | RGB | PBR | RGB | 56.0 | 50.7 | 68.4 | 41.4 | 33.8 | 70.4 | 62.1 | 54.7 | 47.39 |
| 11 | ZeroPose-Multi-Hypo-Refinement [3, 43] | 2023 | DNN | All | SAM + ImageBind | MegaPose | RGB-D | PBR+real | RGB-D | 49.3 | 34.2 | 79.0 | 39.6 | 46.5 | 62.9 | 62.3 | 53.4 | 18.97 |
| 12 | MegaPose-CNOS_fastSAM [33, 43] | 2023 | DNN | All | CNOS-fastSAM | MegaPose | RGB | PBR | RGB | 49.9 | 47.7 | 65.3 | 36.7 | 31.5 | 65.4 | 60.1 | 50.9 | 31.72 |
| 13 | ZeroPose-One-Hypo [3] | 2023 | DNN | All | SAM + ImageBind | MegaPose | RGB-D | PBR+real | RGB-D | 27.2 | 15.6 | 53.6 | 30.7 | 36.2 | 46.2 | 34.1 | 34.8 | 9.76 |
| 14 | GenFlow-coarse | 2023 | DNN | All | CNOS-fastSAM | – | RGB | PBR | RGB | 25.0 | 21.5 | 30.0 | 16.8 | 15.4 | 28.3 | 27.7 | 23.5 | 3.84 |
| # | Method | …based on | Year | Data | …type | APC | Time |
| 1 | GDet2023 | YOLOv8 | 2023 | RGB | PBR+real | 79.8 | .204 |
| 2 | GDRNPPDet | YOLOX | 2022 | RGB | PBR+real | 77.3 | .081 |
| 3 | GDet2023-PBR | YOLOv8 | 2023 | RGB | PBR | 76.9 | .204 |
| 4 | GDRNPPDet | YOLOX | 2022 | RGB | PBR | 73.8 | .081 |
| 5 | Extended_FCOS | FCOS | 2022 | RGB | PBR+real | 72.1 | .030 |
| 6 | Extended_FCOS | FCOS | 2022 | RGB | PBR | 66.7 | .030 |
| 7 | DLZDet | DLZDet | 2022 | RGB | PBR | 65.6 | - |
| 8 | CosyPose | Mask R-CNN | 2020 | RGB | PBR+real | 60.5 | .054 |
| 9 | CosyPose | Mask R-CNN | 2020 | RGB | PBR | 55.7 | .055 |
| 10 | FCOS-CDPN | FCOS | 2022 | RGB | PBR | 50.7 | .047 |
| # | Method | …based on | Year | Data | …type | APC | Time |
| 1 | ZebraPoseSAT | GDRNPP+Zebra | 2023 | RGB | PBR+real | 61.9 | .080 |
| 2 | ZebraPoseSAT | GDRNPP+Zebra | 2022 | RGB | PBR+real | 58.7 | .080 |
| 3 | ZebraPoseSAT | CDPNv2+Zebra | 2023 | RGB | PBR+real | 57.9 | .080 |
| 4 | ZebraPoseSAT | CDPNv2+Zebra | 2022 | RGB | PBR+real | 57.8 | .080 |
| 5 | ZebraPoseSAT | CosyPose+Zebra | 2022 | RGB | PBR | 53.8 | .080 |
| 6 | ZebraPoseSAT | CDPNv2+Zebra | 2022 | RGB | PBR | 52.3 | .080 |
| 7 | DLZDet | DLZDet | 2022 | RGB | PBR+real | 49.6 | - |
| 8 | DLZDet | DLZDet | 2022 | RGB | PBR | 42.9 | - |
| 9 | CosyPose | Mask R-CNN | 2020 | RGB | PBR+real | 40.5 | .054 |
| 10 | CosyPose | Mask R-CNN | 2020 | RGB | PBR | 36.2 | .055 |
| # | Method | Year | Train. im. | …type | Test. im. | APC | Time |
| 1 | CNOS_FastSAM [43] | 2023 | RGB | PBR | RGB | 42.8 | 0.221 |
| 2 | CNOS_SAM [43] | 2023 | RGB | PBR | RGB | 36.1 | 1.847 |
| 3 | ZeroPose [3] | 2023 | RGB | PBR | RGB | 34.1 | 3.821 |